The volume of a rectangular aquarium is 900 gallons. the bottom is 4ft by 6ft. Find the height of the tank. (hint: there are 7.5 gallons per cubic foot)
I just need to know how you would set up the equation?

Answer by ptaylor(2198)

Volume of a rectangular acquarium = Length*Width*Height, or
V=LWH; H=V/(LW)
First, lets find the volume:
Volume=900/(7.5)=120 cu ft
Sooo, our equation to solve is:
H=120/(4*6)=120/24=5ft

Hope this helps---ptaylor